### v1

# Rethinking Data Input for Point Cloud Upsampling

Abstract
In recent years, point cloud upsampling has been widely applied in fields such as 3D reconstruction and surface generation. However, existing point cloud upsampling inputs are all patch based, and there is no research discussing the differences and principles between point cloud model full input and patch based input. In order to compare with patch based point cloud input, this article proposes a new data input method, which divides the full point cloud model to ensure shape integrity while training PU-GCN. This article was validated on the PU1K and ABC datasets, but the results showed that Patch based performance is better than model based full input i.e. Average Segment input. Therefore, this article explores the data input factors and model modules that affect the upsampling results of point clouds.
Keywords:
Deep learning Point cloud Upsampling Data input 3D reconstruction
1 Introduction
The improvement of 3D scanners in 3D sensors, as well as the rapid development of deep learning in computer vision, have led to a gradual shift in the popularity of computer vision technology from 2D image to 3D model. Especially in the direction of autonomous driving and robots that exchange with the physical world, there is a particular focus on three-dimensional vision [5, 21, 17, 28].
Vis-a-vis other representations such as Mesh, point clouds are an effective way to represent three-dimensional shapes that are widely used in the aforementioned fields. In addition, the task of reverse engineering CAD models from 3D point clouds is receiving increasing attention [1]. Reverse engineering involves transforming physical objects into digital models for analysis, visualization, and manufacturing. In the process of redesigning the model obtained from reverse engineering, although there are Application Programming Interface (API) available in CAD software [9], if these designs involve reverse engineering and highly customized and personalized designs, CAD APIs are powerless [22]. Therefore, the generation of model point clouds is crucial.
Figure 1: The original Mesh models have been transformed into a point cloud through patch based sampling and the Average Segment method proposed in this article.
It is beyond dispute that upsampling is a type of generative task. The point cloud upsampling technology based on deep learning has developed with the development of deep learning, extracting multi-resolution features from the initial PU-Net [26] on the basis of pointnet++ [18]. Then, multiple independent multi-layer perceptrons (MLPs) are introduced to generate the required dense point cloud. Moving on to MPU/3PU [25], KNN is used for interpolation and EdgeConv [24] is used to extract features for hierarchical reconstruction through recursive upsampling. Implementing upsampling PU-GAN [12] based on GAN [6] generation architecture, guiding the generator to upsample and approximate the real point cloud. PU-GCN [19] modifies the upsampling module from simple duplicate to Nodeshuffle based on feature extraction in MPU, achieving very good results. And Transformer’s point cloud upsampling network [20, 23]. There are also methods to create coarse point clouds and then refine them by combining global and local methods [13].
The above methods mainly utilize Euclidean spatial features at potential spatial points, upsampling from low resolution to high resolution. The inherent limitations have also been addressed by the work of upsampling point clouds at any scale [27, 8] and capturing local information [10].
Nevertheless, these works still have limitations, which are on the training dataset. Existing point cloud upsampling work [19, 20, 27, 8, 10, 15] without exception used the PU1K dataset [19], which was used based on the methods of MPU and PU-GAN [25, 12]. In order to obtain training data by randomly sampling the entire 3D model using Patch under conditions of limited computing power, especially limited graphics memory. Undoubtedly, Patch based data input can reduce memory consumption and computational complexity, but it can bring about boundary effects and continuity loss in reverse engineering. It is necessary to explore a new point cloud input method to avoid the above issues. Moreover, the existing point cloud upsampling work focuses on 3D models similar to the PU1K dataset, while neglecting the exploration of point cloud upsampling for parts and CAD. In addition to conducting experiments on the PU1K dataset, this article will also discuss the upsampling effect of point cloud models on workpieces on the ABC dataset [11]. In addition, in order to achieve personalized point cloud model generation in one step after reverse engineering, it is necessary to understand which part of the encoder and decoder, namely feature extraction and data expansion modules, determines the performance of upsampling in point cloud upsampling.
Based on the above issues, in order to explore why numerous point cloud sampling data are still based on Patch input, this paper proposes a new data input method. Unlike dividing point clouds according to Patch, this article takes the point cloud model uniformly and randomly divided and corresponds them one by one as input, which is named as Average Segment (AS). After verifying the method proposed in this article, experiments were conducted on point cloud upsampling using the PU-GCN [19] network, and cross validation was performed on both the PU1k and ABC datasets. And during the experiment, the encoder and decoder of PU-GCN [19] were ablated.
In order to solve the above problems, this paper proposes a new data input method. Unlike dividing point clouds according to patches, this paper uniformly and randomly divides point cloud models and uses them as inputs one by one. After verifying the method proposed in this article, experiments were conducted on point cloud upsampling using the PU-GCN [19], and cross validation was performed on both the PU1k [19] and ABC datasets [11]. And during the experiment, the encoder and decoder of PU-GCN [19] were ablated.
The contributions of this article are as follows:
-Proposing a new data input method for point cloud upsampling, namely the Average Segment (AS) method. In theory, the boundary effects and continuity issues brought about by two patch and uniform segmentation methods were discussed.
-Verifying the effectiveness of Patch based method through experiments with AS method and explains the reasons why patch input is superior to AS input.
-Conducting ablation experiments on the main structure of PU-GCN, determining that upsampling mainly relies on the extended feature modules in the decoder, and pointing out the key link in point cloud generation.
2 Methodology
2.1 Overview
As shown in Fig. 1, given a sparse point cloud $\mathcal{P}\in\mathbb{R}^{N\times 3}$, the input methods based on Patch and the AS method proposed in this paper are used as data inputs, with PU-GCN as the network backbone, to generate a dense point cloud $\mathcal{P^{\prime}}\in\mathbb{R}^{rN\times 3}$, where $r$ is the upsampling scale. The reason for choosing PU-GCN as the backbone of the network is that MPU and PU-GCN have similar feature extraction methods. Both use KNN to construct a graph of the input point cloud, with nodes being points and edges defining $k$ nearest point neighborhoods. Then, EdgeConv uses these distances to calculate the edge features of each point. The subsequent PU-Transformer in the upsampling module adopted the Nodeshuffle proposed by PU-GCN, which forms a dense feature map compared to MPU’s Duplicate.
Incidentally, where the input training data of MPU and PU-GAN, and the PU1k dataset proposed by PU-GCN both convert an entire mesh into a point cloud for upsampling and divide the mesh into multiple patches, and then upsample each patch separately. Accordingly, this entire point cloud has its own advantages and disadvantages as two methods for training and sampling into patches.
The disadvantage of converting the entire mesh into a point cloud for upsampling: It may consume a lot of memory resources, especially for complex high-resolution meshes. This may result in insufficient memory or slow processing speed. And for computational complexity, converting large meshes directly into point clouds requires a significant amount of computing resources and time. When performing point cloud upsampling, this high computational complexity will further increase. In addition, the point cloud converted from the entire mesh may be very dense, which may contain a large number of unnecessary redundant points. This brings about the issue of density, as these redundant points increase computational complexity and may lead to overfitting of the model.
Ergo, in order to avoid memory consumption and computational complexity, the previously proposed patch based end-to-end training has emerged. In situations where resources are limited, dividing the mesh into multiple patches for upsampling is more practical and efficient. But there are also disadvantages to this.
The disadvantage of dividing the mesh into multiple patches for upsampling: Frankly, information is missing. Dividing the mesh into multiple patches may result in the loss of important local features, as each patch may not contain complete information about the entire object. This may affect the quality of upsampling. In addition, due to the existence of boundaries between patches, this brings about boundary effects. When dividing the mesh into multiple patches for upsampling, discontinuity or defects may occur. And therefore, the results sampled on each patch may not be continuous enough, as they are processed independently. This may affect the continuity and consistency of the final point cloud.
The AS i.e. Average Segment method proposed in this article converts the entire mesh into a point cloud and performs uniform random sampling and segmentation, which can partially avoid some shortcomings of dividing the mesh into multiple patches. This is because we retained the overall local features in the point cloud through random sampling, and avoided boundary effects and information loss caused by dividing the mesh into discrete segments.
Whereas actually, there are still some potential disadvantages: especially in terms of continuity. Although the overall local features are retained during sampling, point cloud data may not be continuous enough in some areas due to random sampling, and further processing may be needed to ensure the continuity and consistency of the point cloud.
Next, this article will theoretically discuss the differences between the two input methods.
(a) Patch based data input
(b) AS method data input
Figure 2: Schematic diagram of two data input methods for point cloud upsampling
2.2 Data Input Based on Patch
In this case, whichs is showing in Fig. 2(a), it is necessary to divide the Mesh into multiple patches. Assuming the entire Mesh contains $M$ vertices, divide it into $K$ patches. $\mathcal{P}_{k}$ contains $M_{k}$ vertices, where $\sum_{k=1}^{K}M_{k}\geq M$. Convert each patch to a point cloud and perform upsampling. After converting the vertices of each patch to a point cloud, it is assumed that the point cloud contains $N_{k}$ points and we want to upsample them into $N_{k}^{\prime}$ points, typically $N_{k}^{\prime}>N_{k}$.
Process each patch using a graph convolutional network, and record the output of the graph convolutional network as $\mathbf{h}_{i}$represents the feature representation of the $i$-th point. Then, use NodeShuffle technique to generate more points.
Let the graph convolutional network have $L$ layers, and the weight matrix of each layer is $\mathbf{W}_{l}$, then the output of layer $l$is:
$\mathbf{H}^{(l)}=\sigma(\mathbf{A}\mathbf{H}^{(l-1)}\mathbf{W}^{(l)})$
Among them:
--- - $\mathbf{A}$ is an adjacency matrix.
--- - $\mathbf{H}^{(0)}=\mathbf{X}$ representing the feature matrix of the input point.
--- - $\sigma$ is the activation function.
And, we obtain a new feature representation $\mathbf{H}^{(L)}$and use NodeShuffle to generate a new set of points:
$\mathcal{P}_{k}^{\prime}=\text{NodeShuffle}(\mathcal{P}_{k})$
Therefore, there is an upsampled point cloud set, which concatenates all patch upsampled point cloud sets into a complete point cloud:
$\mathcal{P}^{\prime}=\bigcup_{k=1}^{K}\mathcal{P}_{k}^{\prime}$
2.3 Data Input Based on Average Segment
In this process, like the example in Fig. 2, we can assume that the distribution of sampling points follows a uniform distribution. Assuming that the range of a point cloud is a bounded three-dimensional spatial area, we can use uniform distribution to simulate the sampling process of points.
Specifically, assuming the spatial range of a point cloud is $[x_{\text{min}},x_{\text{max}}]\times[y_{\text{min}},y_{\text{max}}]\times[z_{\text{min}},z_{\text{max}}]$, we can consider the coordinates of each point $(x,y,z)$ as three independent random variables, whose distributions are $\text{Uniform}(x_{\text{min}},x_{\text{max}})$, $\text{Uniform}(y_{\text{min}},y_{\text{max}})$ and $\text{Uniform}(z_{\text{min}},z_{\text{max}})$.
After converting the entire mesh to a point cloud, uniform random sampling is performed. Assuming we want to randomly sample $M$ points from $N$ points as a subset, where $M$ is much smaller than $N$. The probability of each point being selected is $p=\frac{M}{N}$.
Assuming we have conducted $N$ independent Bernoulli experiments with a probability of success of $p$, then the number of successful experiments $X$ follows a binomial distribution $B(N,p)$. The selected number of points $S$ follows a binomial distribution $B(N,p)$:
$S=\sum_{i=1}^{N}X_{i},\quad S\sim B(N,p)$
Selected point set $\mathcal{P}_{\text{sampled}}$ can be expressed as:
$\mathcal{P}_{\text{sampled}}=\{\mathbf{p}_{i}\mid X_{i}=1\}$
Once sampling is completed, obtain a point set containing $M$ points $\mathcal{P}_{\text{sampled}}$. We hope to divide it into $K$ subsets, each containing $\frac{M}{K}$ points.
The $k$-th subset is denoted as $\mathcal{P}_{k}$, where $k=1,2,\ldots,K$. Assuming that $M$ can be divided by $K$, the range of points for each subset is:
$\mathcal{P}_{k}=\{\mathbf{p}_{i}\mid(k-1)\cdot\frac{M}{K}+1\leq i\leq k\cdot\frac{M}{K}\}$
Assuming the $k$-th subset $\mathcal{P}_{k}$ contains $N_{k}$ points, and we want to upsample it into a subset $\mathcal{P}_{k}^{\prime}$ containing $N_{k}^{\prime}$ points, where $N_{k}^{\prime}>N_{k}$.
During the upsampling process, use graph convolutional networks to process each subset. Assuming a graph convolutional network has $L$ layers of convolution, and the weight matrix of each layer is $\mathbf{W}_{l}$, then the output of layer $l$ is:
$\mathbf{H}^{(l)}=\sigma(\mathbf{A}\mathbf{H}^{(l-1)}\mathbf{W}^{(l)})$
The pseudocode for the above method is showing in Algorithm 1.
Input: Data matrix $\mathbf{data}$ of shape $(N,P,3)$, number of clusters $K$
#$N\leftarrow\text{number of objects}\leftarrow\mathbf{data}.\text{shape}[0]$
#$P\leftarrow\text{number of points}\leftarrow\mathbf{data}.\text{shape}[1]$
1
[…]
3for each segment in segmented_data do
4      $\text{segment}\leftarrow\text{reshape}(\text{segment},(-1,\frac{P}{K},3))$
5 end for
Algorithm 1 Point Cloud Upsampling with Patch Segmentation
3 Experiments
3.1 Settings
Datasets: In general speaking, the experiment in this article applied the following two 3D benchmark datasets:
  • PU1K: This is a point cloud upsampling dataset introduced in PU-GCN [19]. Overall, PU1K consists of 1147 3D models, divided into 1020 training samples and 127 testing samples. The training set includes 120 3D models from the PU-GAN dataset, as well as 900 different models collected from ShapeNetCore [4]. The test set includes 27 models from PU-GAN and over 100 models from ShapeNetCore. Covering 50 object categories.
The original PU1K dataset was used to adapt to patch based upsampling pipelines, and the training data was generated from patches in the 3D mesh through Poisson sampling. Specifically, the training data includes 69000 samples, each with low resolution: 256 input points and high resolution: 1024 groundtruth points which is four times larger than input. In addition, the method based on AS mentioned in Chapter 2 above also modified the PU1K dataset by uniformly dividing the original 1020 training samples into 8 sub-models that maintain the shape integrity of the model from 2048 points as input in a single model and 8192 points in ground truth. The training data consists of 8160 samples. Similarly, each sample has low resolution: 256 input points and high resolution: 1024 groundtruth points.
  • ABC Dataset: The ABC Dataset [11] is a dataset containing one million computer-aided design (CAD) models, aimed at promoting research on geometric deep learning methods and applications. It is currently one of the largest publicly available datasets for geometric deep learning research.
The vast CAD dataset brings diversity, with models covering a wide range of geometric shapes and categories, including architecture, furniture, mechanical parts, and more. And each model is composed of explicitly parameterized curves and surfaces, which can provide real data in areas such as differentiation, crack segmentation, geometric feature detection, and shape reconstruction. This article is based on a subset of ABC 10k in the ABC dataset, which is a CAD model with a quantity of 10000. On this basis, two 4x upsampling experiments were conducted, with the inputs being the complete model and the patch model, respectively. Each sample has low resolution: 512 input points and high resolution: 2048 groundtruth points.
Training details: Straightforwardly, the PU-GCN used in this article, where the network structure is shown in the architecture of (a) in Fig. 3, is implemented using PyTorch. Due to TensorFlow and CUDA versions, this article is unable to re-create the original author’s Python environment in PU-GCN within the TensorFlow deep learning framework. And the experiments were conducted on a single NVIDIA GeForce RTX 3060 LHR GPU running on a Ubuntu 22.04 operating system. In terms of training hyperparameters, this article follows the settings of the original PU-GCN in terms of network structure. Notwithstanding, a batch size of $32$ with $50$ training epochs was used, with an initial learning rate of $5\times 10^{-4}$, a decay rate of $0.05$, and so on. In addition, the loss function also uses Chamfer Distance loss to train PU-GCN, minimizing the average nearest point distance between the input set $\mathcal{P}\in\mathbb{R}^{N\times 3}$ and the output set $\mathcal{P^{\prime}}\in\mathbb{R}^{rN\times 3}$, thus achieving efficient and effective convergence.
Figure 3: The PU-GCN network structure used in the experiment is shown in the figure, where (a) is the original PU-GCN structure. (b), (c) and (d) are all architectures used in the ablation experiment to explore the modules that have the greatest impact on upsampling.
3.2 Evaluation Metrics
During the testing process, this article adopted the common practices used in previous work on point cloud upsampling [19, 20, 27, 8, 10, 15]. That is to say, this article will cut the input point cloud into multiple subset Patch and AS of all N points. Then, using the PU-GCN model to upsample the subset Patch and AS at a scale of $r$.
In the 4 times upsampling experiment of this article, three types of low resolution to high-resolution sampling were performed. Each test sample has a low resolution point cloud consisting of 256, 512 and 1048 points, and a high-resolution point cloud consisting of 1024,2048 and 8196 points. This article quantitatively evaluates the upsampling performance of PU-GCN under two methods based on three widely used indicators: (i) Chamfer distance (CD) [7], (ii) Hausdorff distance (HD) [2], and (iii) model parameter quantity. The lower the values under these indicators, the better the upsampling performance.
3.3 Results
PU1K dataset: Tab. 1 shows the quantitative results of using patch method and uniform segmentation method with PU-GCN network training on the PU1K dataset. By upsampling at resolutions of 256-1024 and 2048-8192, it can be seen that the Patch based method outperforms the AS method proposed in this paper in all two metrics of distance measurement.
Table 1: The data input based on the Patch method and AS method is obtained using PU-GCN in PU1K.
Datasets Test Scale Method Results (×$(\times 10^{-3})$
CD $\downarrow$ HD $\downarrow$
PU1K $256\to 1024$ $256\to 1024$ Patch 35.459 54.884
AS 35.529 54.923
$512\to 2048$ Patch 25.689 41.990
AS 25.616 42.344
$2048\to 8192$ Patch 14.751 23.915
AS 14.961 27.822
ABC10K dataset: we also conducted point cloud upsampling experiments using the ABC10K dataset. Perform upsampling at a resolution of 512-2048. As shown in Tab. 2, in the 4$\times$ upsampling experiment, the two evaluation metrics of the Patch based method outperformed the AS method proposed in this paper in all two metrics of distance measurement.
Table 2: The data input based on the Patch method and AS method is obtained using PU-GCN in ABC10K.
Datasets Test Scale Method Results (×$(\times 10^{-3})$
CD $\downarrow$ HD $\downarrow$
ABC10K $512\to 2048$ $512\to 2048$ Patch 42.861 84.128
AS 47.991 95.154
Although the idea, which is originated from overall uniform random sampling, article proposed name Average Segment (AS) method has many advantages in theory, actual experimental results show that the Patch based upsampling method performs better on Chamfer distance (CD) and Hausdorff distance (HD). One possible reason for this may be that when upsampling on a patch, each patch contains fewer point clouds, and the graph convolutional network can extract local features more finely. Due to each patch of the point cloud being within a small range, local features are easier to capture and reconstruct, resulting in excellent performance on CD and HD. Other possibilities that could cause this situation will be discussed in Chapter 4.
3.4 Ablations
Test sets cross validation: This article will conduct cross validation on the PU-GCN trained on the PU1K and ABC10K datasets under the validation sets of both datasets. To understand the impact of the model used for training data on upsampling results. The results are shown in Tab. 3. When the Patch method and AS method are trained on the PU1K dataset and tested on the ABC10K dataset, the CD and HD metrics are significantly superior to those trained and tested on the ABC10K dataset. However, when trained on the ABC10K dataset and tested on the PU1K dataset, there was no significant difference between the CD and HD indicators. It should be noted that in the aforementioned cross validation, the Patch method is still better than the AS method.
Table 3: The results of swapping test sets based on Patch method and AS method on two datasets.
Training Sets Method Test Sets Results (×$(\times 10^{-3})$
CD $\downarrow$ HD $\downarrow$
PU1K $256\to 1024$ Patch ABC10K $512\to 2048$ 22.300 43.386
AS 41.286 82.695
ABC10K $512\to 2048$ Patch PU1K $256\to 1024$ 35.260 53.203
AS 38.494 57.212
Patch PU1K $512\to 2048$ 25.200 39.345
AS 29.225 43.932
Patch PU1K $2048\to 8192$ 13.121 21.707
AS 18.689 27.688
Architecture Impact: This article breaks down the important modules of PU-GCN, aiming to explore which part has the greatest impact on point cloud upsampling. To this end, the original network backbone has been increased or decreased in three ways: removing DenseGCN (b), adding Refiner (c), and removing DenseGCN while adding Refiner (d), all of them are showed in Fig. 3. The experiment was trained on the PU1K dataset and tested on the test sets of the PU1K dataset and ABC10K dataset. The results are shown in Tab. 4. The results showed that upsampling on the test set with three resolutions of 256-1024, 512-2048 and 2048-8192 in PU1K performed better than the original PU-GCN with DenseGCN, regardless of the presence or absence of Refiner, under the condition of removing DenseGCN. Under the upsampling scale of 2048-8192, the uniform segmentation method that removes DenseGCN is better than the Patch method.
Table 4: Perform ablation experiments on different architectures of PU-GCN using two data input methods.
Network Backbone Method Test Sets Results (×$(\times 10^{-3})$ Model size (MB)
CD $\downarrow$ HD $\downarrow$
PU-GCN Patch PU1K $256\to 1024$ 35.459 54.884 66.246
AS 35.529 54.923
Patch PU1K $2048\to 8192$ 14.751 23.915
AS 14.961 27.822
without Inception DenseGCN Patch PU1K $256\to 1024$ 35.061 54.340 99.152
AS 35.229 52.819
Patch PU1K $2048\to 8192$ 13.736 22.573
AS 13.504 23.130
with Refiner Patch PU1K $256\to 1024$ 35.075 54.227 87.773
AS 35.860 53.652
Patch PU1K $2048\to 8192$ 14.361 23.275
AS 14.817 26.860
with Refiner but without Inception DenseGCN Patch PU1K $256\to 1024$ 34.309 54.070 130.945
AS 35.158 53.434
Patch PU1K $2048\to 8192$ 13.721 22.245
AS 13.441 22.899
3.5 Visualizations
Visualize the above experimental results now, which directly displays qualitative comparisons. The qualitative results of upsampling using different point cloud datasets are shown in Fig. 4 and Fig. 5. It can be seen that the input based on Patch performs better in distinguishing edges on the boundaries compared to the AS method. This is particularly evident on the CAD based ABC dataset, as shown in Fig. 5. However, in Fig. 4, it can be seen that the AS method performs better locally in the model. Relatively speaking, it is denser and the connections between points are tighter. In addition, even though DenseGCN performs better on CD and HD on the PU1K dataset, this results in the point cloud lacking local features in visualization. On the other hand, Refiner can solve the problem of identifying local features in the model, achieving the utilization of both global and local features.
Figure 4: Visualization of PU-GCN with Different Architectures on the PU1K Dataset.
Figure 5: Visualization example on the ABC dataset.
4 Discussion and Future Work
4.1 Discussion
Now, we will focus on why the patch upsampling method performs better in PU-GCN experiments. Although the AS random sampling method proposed in this article has many advantages, the actual experimental results are not as good as the input method based on Patch in terms of Chamfer distance (CD) and Hausdorff distance (HD). This phenomenon can be explained by the following points:
More effective local feature extraction: When upsampling on AS method, each AS slice contains fewer point clouds, and the graph convolutional network can extract local features more finely. Due to each AS slice of the point cloud being within a small range, local features are easier to capture and reconstruct, resulting in excellent performance on CD and HD.
Set $\mathcal{P}_{k}$ represents the $k$-th AS slice, $\mathcal{P}_{k}=\{\mathbf{p}_{i}\mid i\in I_{k}\}$, where $I_{k}$ is the point index set for AS $\mathcal{P}_{k}$. For each AS slice $\mathcal{P}_{k}$, its point cloud can be represented as $\mathbf{X}_{k}\in\mathbb{R}^{N_{k}\times d}$, where $N_{k}$ is the number of AS slices and $d$ is the dimension of the points.
In a graph convolutional network, the output of the $l$-th layer is represented as:
$\mathbf{H}_{k}^{(l)}=\sigma(\mathbf{A}_{k}\mathbf{H}_{k}^{(l-1)}\mathbf{W}^{(l)})$
In an AS slice, the adjacency matrix $\mathbf{A}_{k}$ is denser, with tighter connections between points, making convolution operations more effective in capturing local features.
Local consistency of AS slices: When dealing with smaller AS slices, the local consistency of point clouds is stronger, that is, the similarity between points is higher. This enables graph convolutional networks to better utilize these local consistency for learning and reconstruction, reducing local errors introduced by Patch based method.
The AS method deals with point clouds with strong local consistency, which can be measured by the local covariance matrix.
For AS slice $\mathcal{P}_{k}$, its point cloud $\mathbf{X}_{k}$ have the local covariance matrix is:
$\mathbf{C}_{k}=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}(\mathbf{x}_{i}-\mathbf{\mu}_{k})(\mathbf{x}_{i}-\mathbf{\mu}_{k})^{\top}$
Where $\mathbf{\mu}_{k}$ is mean vector of midpoint of AS slice $\mathcal{P}_{k}$ :
$\mathbf{\mu}_{k}=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}\mathbf{x}_{i}$
Local covariance matrix $\mathbf{C}_{k}$ can measure the local consistency of point clouds. In the case of strong local consistency, the smaller feature values of $\mathbf{C}_{k}$ make local feature extraction more effective.
The effectiveness of boundary processing strategies: Although theoretically AS sampling followed by segmentation can avoid boundary problems, in practical applications, the overlapping and smoothing of boundaries in patch segmentation may be more effective than those after AS sampling, thereby improving the overall quality of upsampling.
Set two adjacent patches as $\mathcal{P}_{k}$ and $\mathcal{P}_{m}$. The overlapping area between them is $\mathcal{O}_{k,m}$, whose point cloud is represented as $\mathbf{X}_{k,m}\in\mathbb{R}^{N_{k,m}\times d}$, where $N_{k,m}$ is the number of overlapping regions.
Smoothing in overlapping areas can reduce errors caused by boundary effects:
$\mathbf{X}_{k,m}^{\prime}=\alpha\mathbf{X}_{k,m}^{(k)}+(1-\alpha)\mathbf{X}_{k,m}^{(m)}$
Among them:
  • $\mathbf{X}_{k,m}^{(k)}$ and $\mathbf{X}_{k,m}^{(m)}$ respectively represent overlapping area point clouds in Patch $\mathcal{P}_{k}$ and $\mathcal{P}_{m}$.
  • $\alpha\in[0,1]$ is the smoothing coefficient
The smoothing of this overlapping area can effectively reduce boundary effects and improve the overall quality of sampling.
Insufficient utilization of global information: Although AS methods retain global information, some global information may be lost during processing and segmentation, resulting in a decrease in the quality of the upsampling point cloud. On the contrary, although the patch method processes local information, it can still reconstruct global information well through the synthesis of multiple patches.
The method for AS as a whole is $\mathcal{P}_{\text{sampled}}=\{\mathbf{p}_{i}\mid i\in I_{\text{sampled}}\}$, divide it into $K$ subsets:
$\mathcal{P}_{k}=\{\mathbf{p}_{i}\mid i\in I_{k},I_{k}\subset I_{\text{sampled}},|I_{k}|=\frac{M}{K}\}$
Due to the segmented subset $\mathcal{P}_{k}$ lack of local consistency may lead to a decrease in the quality of feature extraction and reconstruction.
4.2 Future Work
In point cloud 3D model sampling, such as Pu-Edgeformr++ [10], it can achieve good point cloud edge recognition and sampling. However, taking the PU-GCN model used in this article as an example, the existing point cloud upsampling models are not convincing in their performance on the ABC dataset. The method of reverse engineering CAD models follows a series of typical steps: data capture and preprocessing, point cloud segmentation, face classification, and analysis of faces generated by segmentation methods [3]. The poor quantitative and qualitative performance on the CAD dataset ABC may be some challenges due to the heterogeneous and unstructured nature of manufacturing data [16]. In the future, the application of point cloud upsampling in reverse engineering and CAD datasets may require a similar approach to Point2CAD [14], which involves dividing point clouds into clusters corresponding to CAD model faces, fitting each cluster with an analytical surface primitive, and proposing a novel fitting scheme for free-form surfaces based on neural representation.
5 Conclusion
In order to ensure the model integrity of point cloud upsampling, this article mainly studies the input of point cloud data. Specifically, this article proposes a new data input method, also known as a new training model method, called Average Segment. And experiments were conducted on a stable and lightweight PU-GCN model. The traditional patch method still performs slightly better in both CD and HD indicators, but through relevant ablation research and visualization, we can see that the AS method can maintain local features within the model. In future work, we consider combining point cloud models with mesh face clusters to perform pairwise surface intersections to identify object edges, thereby achieving better edge, local, and global representations.

### v2

@@
 Rethinking Data Input for Point Cloud Upsampling
-Abstract
-In recent years, point cloud upsampling has been widely applied in fields such as 3D reconstruction and surface generation.
-However, existing point cloud upsampling inputs are all patch based, and there is no research discussing the differences and principles between point cloud model full input and patch based input.
-In order to compare with patch based point cloud input, this article proposes a new data input method, which divides the full point cloud model to ensure shape integrity while training PU-GCN.
-This article was validated on the PU1K and ABC datasets, but the results showed that Patch based performance is better than model based full input i.e.
+Abstract.
+Point cloud upsampling is crucial for tasks like 3D reconstruction.
+While existing methods rely on patch-based inputs, and there is no research discussing the differences and principles between point cloud model full input and patch based input.
+Ergo, we propose a novel approach using whole model inputs i.e.
 Average Segment input.
-Therefore, this article explores the data input factors and model modules that affect the upsampling results of point clouds.
-Keywords:
-Deep learning Point cloud Upsampling Data input 3D reconstruction
-1 Introduction
+Our experiments on PU1K and ABC datasets reveal that patch-based inputs consistently outperform whole model inputs.
+To understand this, we will delve into factors in feature extraction, and network architecture that influence upsampling results.
+Deep learning, Point cloud, Upsampling, Data input, 3D reconstruction
+ccs: Computing methodologies Computer vision
+1. Introduction
 The improvement of 3D scanners in 3D sensors, as well as the rapid development of deep learning in computer vision, have led to a gradual shift in the popularity of computer vision technology from 2D image to 3D model.
-Especially in the direction of autonomous driving and robots that exchange with the physical world, there is a particular focus on three-dimensional vision [5, 21, 17, 28].
+Especially in the direction of autonomous driving and robots that exchange with the physical world, there is a particular focus on three-dimensional vision (Endres et al., 2013; Roriz et al., 2021; Naseer et al., 2018; Zhou and Tuzel, 2018).
+Figure 1. The original Mesh models have been transformed into a point cloud through patch based sampling and the Average Segment method proposed in this article.
 Vis-a-vis other representations such as Mesh, point clouds are an effective way to represent three-dimensional shapes that are widely used in the aforementioned fields.
-In addition, the task of reverse engineering CAD models from 3D point clouds is receiving increasing attention [1].
-Reverse engineering involves transforming physical objects into digital models for analysis, visualization, and manufacturing.
-In the process of redesigning the model obtained from reverse engineering, although there are Application Programming Interface (API) available in CAD software [9], if these designs involve reverse engineering and highly customized and personalized designs, CAD APIs are powerless [22].
+In addition, the task of reverse engineering CAD models from 3D point clouds is receiving increasing attention (Alai, 2013).
+In the process of redesigning the model obtained from reverse engineering, although there are Application Programming Interface (API) available in CAD software (Kao and Lin, 1998), if these designs involve reverse engineering and highly customized and personalized designs, CAD APIs are powerless (Shatnawi et al., 2017).
 Therefore, the generation of model point clouds is crucial.
-Figure 1: The original Mesh models have been transformed into a point cloud through patch based sampling and the Average Segment method proposed in this article.
-It is beyond dispute that upsampling is a type of generative task.
-The point cloud upsampling technology based on deep learning has developed with the development of deep learning, extracting multi-resolution features from the initial PU-Net [26] on the basis of pointnet++ [18].
+The point cloud upsampling technology based on deep learning has developed with the development of deep learning, extracting multi-resolution features from the initial PU-Net (Yu et al., 2018) on the basis of pointnet++ (Qi et al., 2017).
 Then, multiple independent multi-layer perceptrons (MLPs) are introduced to generate the required dense point cloud.
-Moving on to MPU/3PU [25], KNN is used for interpolation and EdgeConv [24] is used to extract features for hierarchical reconstruction through recursive upsampling.
-Implementing upsampling PU-GAN [12] based on GAN [6] generation architecture, guiding the generator to upsample and approximate the real point cloud.
-PU-GCN [19] modifies the upsampling module from simple duplicate to Nodeshuffle based on feature extraction in MPU, achieving very good results.
-And Transformer’s point cloud upsampling network [20, 23].
-There are also methods to create coarse point clouds and then refine them by combining global and local methods [13].
+Moving on to MPU/3PU (Yifan et al., 2019), KNN is used for interpolation and EdgeConv (Wang et al., 2019) is used to extract features for hierarchical reconstruction through recursive upsampling.
+Implementing upsampling PU-GAN (Li et al., 2019) based on GAN (Goodfellow et al., 2020) generation architecture, guiding the generator to upsample and approximate the real point cloud.
+PU-GCN (Qian et al., 2021) modifies the upsampling module from simple duplicate to Nodeshuffle based on feature extraction in MPU, achieving very good results.
+And Transformer’s point cloud upsampling network (Qiu et al., 2022; Vaswani et al., 2017).
+There are also methods to create coarse point clouds and then refine them by combining global and local methods (Li et al., 2021).
 The above methods mainly utilize Euclidean spatial features at potential spatial points, upsampling from low resolution to high resolution.
-The inherent limitations have also been addressed by the work of upsampling point clouds at any scale [27, 8] and capturing local information [10].
+The inherent limitations have also been addressed by the work of upsampling point clouds at any scale (Zhao et al., 2022; He et al., 2023) and capturing local information (Kim et al., 2024).
 Nevertheless, these works still have limitations, which are on the training dataset.
-Existing point cloud upsampling work [19, 20, 27, 8, 10, 15] without exception used the PU1K dataset [19], which was used based on the methods of MPU and PU-GAN [25, 12].
-In order to obtain training data by randomly sampling the entire 3D model using Patch under conditions of limited computing power, especially limited graphics memory.
+Existing point cloud upsampling work (Qian et al., 2021; Qiu et al., 2022; Zhao et al., 2022; He et al., 2023; Kim et al., 2024; Mao et al., 2022) without exception used the PU1K dataset (Qian et al., 2021), which was used based on the methods of MPU and PU-GAN (Yifan et al., 2019; Li et al., 2019).
+To obtain training data by randomly sampling the entire 3D model using Patch under conditions of limited computing power, especially limited graphics memory.
 Undoubtedly, Patch based data input can reduce memory consumption and computational complexity, but it can bring about boundary effects and continuity loss in reverse engineering.
@@
 Moreover, the existing point cloud upsampling work focuses on 3D models similar to the PU1K dataset, while neglecting the exploration of point cloud upsampling for parts and CAD.
-In addition to conducting experiments on the PU1K dataset, this article will also discuss the upsampling effect of point cloud models on workpieces on the ABC dataset [11].
-In addition, in order to achieve personalized point cloud model generation in one step after reverse engineering, it is necessary to understand which part of the encoder and decoder, namely feature extraction and data expansion modules, determines the performance of upsampling in point cloud upsampling.
+In addition to conducting experiments on the PU1K dataset, this article will also discuss the upsampling effect of point cloud models on workpieces on the ABC dataset (Koch et al., 2019).
 Based on the above issues, in order to explore why numerous point cloud sampling data are still based on Patch input, this paper proposes a new data input method.
 Unlike dividing point clouds according to Patch, this article takes the point cloud model uniformly and randomly divided and corresponds them one by one as input, which is named as Average Segment (AS).
-After verifying the method proposed in this article, experiments were conducted on point cloud upsampling using the PU-GCN [19] network, and cross validation was performed on both the PU1k and ABC datasets.
-And during the experiment, the encoder and decoder of PU-GCN [19] were ablated.
-In order to solve the above problems, this paper proposes a new data input method.
+After verifying the method proposed in this article, experiments were conducted on point cloud upsampling using the PU-GCN (Qian et al., 2021) network, and cross validation was performed on both the PU1k and ABC datasets.
+And during the experiment, the encoder and decoder of PU-GCN (Qian et al., 2021) were ablated.
+To solve the above problems, this paper proposes a new data input method.
 Unlike dividing point clouds according to patches, this paper uniformly and randomly divides point cloud models and uses them as inputs one by one.
-After verifying the method proposed in this article, experiments were conducted on point cloud upsampling using the PU-GCN [19], and cross validation was performed on both the PU1k [19] and ABC datasets [11].
-And during the experiment, the encoder and decoder of PU-GCN [19] were ablated.
+After verifying the method proposed in this article, experiments were conducted on point cloud upsampling using the PU-GCN (Qian et al., 2021), and cross validation was performed on both the PU1k (Qian et al., 2021) and ABC datasets (Koch et al., 2019).
+And during the experiment, the encoder and decoder of PU-GCN (Qian et al., 2021) were ablated.
 The contributions of this article are as follows:
@@
 -Conducting ablation experiments on the main structure of PU-GCN, determining that upsampling mainly relies on the extended feature modules in the decoder, and pointing out the key link in point cloud generation.
-2 Methodology
-2.1 Overview
-As shown in Fig.
-1, given a sparse point cloud $\mathcal{P}\in\mathbb{R}^{N\times 3}$, the input methods based on Patch and the AS method proposed in this paper are used as data inputs, with PU-GCN as the network backbone, to generate a dense point cloud $\mathcal{P^{\prime}}\in\mathbb{R}^{rN\times 3}$, where $r$ is the upsampling scale.
+2. Methodology
+2.1. Overview
+As shown in Figure  1, given a sparse point cloud $\mathcal{P}\in\mathbb{R}^{N\times 3}$, the input methods based on Patch and the AS method proposed in this paper are used as data inputs, with PU-GCN as the network backbone, to generate a dense point cloud $\mathcal{P^{\prime}}\in\mathbb{R}^{rN\times 3}$, where $r$ is the upsampling scale.
 The reason for choosing PU-GCN as the backbone of the network is that MPU and PU-GCN have similar feature extraction methods.
@@
 This brings about the issue of density, as these redundant points increase computational complexity and may lead to overfitting of the model.
-Ergo, in order to avoid memory consumption and computational complexity, the previously proposed patch based end-to-end training has emerged.
-In situations where resources are limited, dividing the mesh into multiple patches for upsampling is more practical and efficient.
+Ergo, to avoid memory consumption and computational complexity, the previously proposed patch based end-to-end training has emerged.
+Dividing the mesh into multiple patches for upsampling is more practical and efficient.
 But there are also disadvantages to this.
@@
 Although the overall local features are retained during sampling, point cloud data may not be continuous enough in some areas due to random sampling, and further processing may be needed to ensure the continuity and consistency of the point cloud.
-Next, this article will theoretically discuss the differences between the two input methods.
-(a) Patch based data input
-(b) AS method data input
-Figure 2: Schematic diagram of two data input methods for point cloud upsampling
-2.2 Data Input Based on Patch
-In this case, whichs is showing in Fig.
-2(a), it is necessary to divide the Mesh into multiple patches.
+2.2. Data Input Based on Patch
+In this case, it is necessary to divide the Mesh into multiple patches.
 Assuming the entire Mesh contains $M$ vertices, divide it into $K$ patches.
@@
 $\mathbf{H}^{(l)}=\sigma(\mathbf{A}\mathbf{H}^{(l-1)}\mathbf{W}^{(l)})$
-Among them:
+where:
 --- - $\mathbf{A}$ is an adjacency matrix.
@@
 $\mathcal{P}^{\prime}=\bigcup_{k=1}^{K}\mathcal{P}_{k}^{\prime}$
-2.3 Data Input Based on Average Segment
-In this process, like the example in Fig.
-2, we can assume that the distribution of sampling points follows a uniform distribution.
+2.3. Data Input Based on Average Segment
+We can assume that the distribution of sampling points follows a uniform distribution.
 Assuming that the range of a point cloud is a bounded three-dimensional spatial area, we can use uniform distribution to simulate the sampling process of points.
-Specifically, assuming the spatial range of a point cloud is $[x_{\text{min}},x_{\text{max}}]\times[y_{\text{min}},y_{\text{max}}]\times[z_{\text{min}},z_{\text{max}}]$, we can consider the coordinates of each point $(x,y,z)$ as three independent random variables, whose distributions are $\text{Uniform}(x_{\text{min}},x_{\text{max}})$, $\text{Uniform}(y_{\text{min}},y_{\text{max}})$ and $\text{Uniform}(z_{\text{min}},z_{\text{max}})$.
 After converting the entire mesh to a point cloud, uniform random sampling is performed.
@@
 $\mathbf{H}^{(l)}=\sigma(\mathbf{A}\mathbf{H}^{(l-1)}\mathbf{W}^{(l)})$
-The pseudocode for the above method is showing in Algorithm 1.
+The pseudocode for the above method is showing in Algorithm  1.
 Input: Data matrix $\mathbf{data}$ of shape $(N,P,3)$, number of clusters $K$
@@
 Algorithm 1 Point Cloud Upsampling with Patch Segmentation
-3 Experiments
-3.1 Settings
+3. Experiments
+3.1. Settings
 Datasets: In general speaking, the experiment in this article applied the following two 3D benchmark datasets:
-  • PU1K: This is a point cloud upsampling dataset introduced in PU-GCN [19].
-Overall, PU1K consists of 1147 3D models, divided into 1020 training samples and 127 testing samples.
-The training set includes 120 3D models from the PU-GAN dataset, as well as 900 different models collected from ShapeNetCore [4].
+  • PU1K: This is a point cloud upsampling dataset introduced in PU-GCN (Qian et al., 2021).
+Overall, PU1K consists of 1147 3D models, including 120 3D models from the PU-GAN dataset, as well as 900 different models collected from ShapeNetCore (Chang et al., 2015).
 The test set includes 27 models from PU-GAN and over 100 models from ShapeNetCore.
 Covering 50 object categories.
-The original PU1K dataset was used to adapt to patch based upsampling pipelines, and the training data was generated from patches in the 3D mesh through Poisson sampling.
-Specifically, the training data includes 69000 samples, each with low resolution: 256 input points and high resolution: 1024 groundtruth points which is four times larger than input.
-In addition, the method based on AS mentioned in Chapter 2 above also modified the PU1K dataset by uniformly dividing the original 1020 training samples into 8 sub-models that maintain the shape integrity of the model from 2048 points as input in a single model and 8192 points in ground truth.
-The training data consists of 8160 samples.
-Similarly, each sample has low resolution: 256 input points and high resolution: 1024 groundtruth points.
-  • ABC Dataset: The ABC Dataset [11] is a dataset containing one million computer-aided design (CAD) models, aimed at promoting research on geometric deep learning methods and applications.
+  • ABC Dataset: The ABC Dataset (Koch et al., 2019) is a dataset containing one million computer-aided design (CAD) models, aimed at promoting research on geometric deep learning methods and applications.
 It is currently one of the largest publicly available datasets for geometric deep learning research.
-The vast CAD dataset brings diversity, with models covering a wide range of geometric shapes and categories, including architecture, furniture, mechanical parts, and more.
-And each model is composed of explicitly parameterized curves and surfaces, which can provide real data in areas such as differentiation, crack segmentation, geometric feature detection, and shape reconstruction.
-This article is based on a subset of ABC 10k in the ABC dataset, which is a CAD model with a quantity of 10000.
-On this basis, two 4x upsampling experiments were conducted, with the inputs being the complete model and the patch model, respectively.
-Each sample has low resolution: 512 input points and high resolution: 2048 groundtruth points.
-Training details: Straightforwardly, the PU-GCN used in this article, where the network structure is shown in the architecture of (a) in Fig.
-3, is implemented using PyTorch.
+Training details: Straightforwardly, the PU-GCN used in this article, where the network structure is shown in the architecture of (a) in Figure  2, is implemented using PyTorch.
 Due to TensorFlow and CUDA versions, this article is unable to re-create the original author’s Python environment in PU-GCN within the TensorFlow deep learning framework.
@@
 Notwithstanding, a batch size of $32$ with $50$ training epochs was used, with an initial learning rate of $5\times 10^{-4}$, a decay rate of $0.05$, and so on.
-In addition, the loss function also uses Chamfer Distance loss to train PU-GCN, minimizing the average nearest point distance between the input set $\mathcal{P}\in\mathbb{R}^{N\times 3}$ and the output set $\mathcal{P^{\prime}}\in\mathbb{R}^{rN\times 3}$, thus achieving efficient and effective convergence.
-Figure 3: The PU-GCN network structure used in the experiment is shown in the figure, where (a) is the original PU-GCN structure.
+Figure 2.
+The PU-GCN network structure used in the experiment is shown in the figure, where (a) is the original PU-GCN structure.
 (b), (c) and (d) are all architectures used in the ablation experiment to explore the modules that have the greatest impact on upsampling.
-3.2 Evaluation Metrics
-During the testing process, this article adopted the common practices used in previous work on point cloud upsampling [19, 20, 27, 8, 10, 15].
+3.2. Evaluation Metrics
+During the testing process, this article adopted the common practices used in previous work on point cloud upsampling (Qian et al., 2021; Qiu et al., 2022; Zhao et al., 2022; He et al., 2023; Kim et al., 2024; Mao et al., 2022).
 That is to say, this article will cut the input point cloud into multiple subset Patch and AS of all N points.
-Then, using the PU-GCN model to upsample the subset Patch and AS at a scale of $r$.
 In the 4 times upsampling experiment of this article, three types of low resolution to high-resolution sampling were performed.
 Each test sample has a low resolution point cloud consisting of 256, 512 and 1048 points, and a high-resolution point cloud consisting of 1024,2048 and 8196 points.
-This article quantitatively evaluates the upsampling performance of PU-GCN under two methods based on three widely used indicators: (i) Chamfer distance (CD) [7], (ii) Hausdorff distance (HD) [2], and (iii) model parameter quantity.
-The lower the values under these indicators, the better the upsampling performance.
-3.3 Results
-PU1K dataset: Tab.
-1 shows the quantitative results of using patch method and uniform segmentation method with PU-GCN network training on the PU1K dataset.
+This article quantitatively evaluates the upsampling performance of PU-GCN under two methods based on three widely used indicators: (i) Chamfer distance (CD) (Hajdu et al., 2012), (ii) Hausdorff distance (HD) (Berger et al., 2013), and (iii) model parameter quantity.
+3.3. Results
+PU1K dataset: Table  1 shows the quantitative results of using patch method and uniform segmentation method with PU-GCN network training on the PU1K dataset.
 By upsampling at resolutions of 256-1024 and 2048-8192, it can be seen that the Patch based method outperforms the AS method proposed in this paper in all two metrics of distance measurement.
-Table 1: The data input based on the Patch method and AS method is obtained using PU-GCN in PU1K.
+Table 1. The data input based on the Patch method and AS method is obtained using PU-GCN in PU1K.
 Datasets Test Scale Method Results (×$(\times 10^{-3})$
@@
 Perform upsampling at a resolution of 512-2048.
-As shown in Tab.
-2, in the 4$\times$ upsampling experiment, the two evaluation metrics of the Patch based method outperformed the AS method proposed in this paper in all two metrics of distance measurement.
-Table 2: The data input based on the Patch method and AS method is obtained using PU-GCN in ABC10K.
+As shown in Table  2, in the 4$\times$ upsampling experiment, the two evaluation metrics of the Patch based method outperformed the AS method proposed in this paper in all two metrics of distance measurement.
+Table 2. The data input based on the Patch method and AS method is obtained using PU-GCN in ABC10K.
 Datasets Test Scale Method Results (×$(\times 10^{-3})$
@@
 AS 47.991 95.154
-Although the idea, which is originated from overall uniform random sampling, article proposed name Average Segment (AS) method has many advantages in theory, actual experimental results show that the Patch based upsampling method performs better on Chamfer distance (CD) and Hausdorff distance (HD).
+Although article proposed name Average Segment (AS) method has many advantages in theory, experimental results show that the Patch based upsampling method performs better on Chamfer distance and Hausdorff distance.
 One possible reason for this may be that when upsampling on a patch, each patch contains fewer point clouds, and the graph convolutional network can extract local features more finely.
-Due to each patch of the point cloud being within a small range, local features are easier to capture and reconstruct, resulting in excellent performance on CD and HD.
-Other possibilities that could cause this situation will be discussed in Chapter 4.
-3.4 Ablations
+Ergo, local features are easier to capture and reconstruct, resulting in excellent performance on CD and HD.
+3.4. Ablations
 Test sets cross validation: This article will conduct cross validation on the PU-GCN trained on the PU1K and ABC10K datasets under the validation sets of both datasets.
 To understand the impact of the model used for training data on upsampling results.
-The results are shown in Tab.
-3.
+The results are shown in Table  3.
 When the Patch method and AS method are trained on the PU1K dataset and tested on the ABC10K dataset, the CD and HD metrics are significantly superior to those trained and tested on the ABC10K dataset.
 However, when trained on the ABC10K dataset and tested on the PU1K dataset, there was no significant difference between the CD and HD indicators.
-It should be noted that in the aforementioned cross validation, the Patch method is still better than the AS method.
-Table 3: The results of swapping test sets based on Patch method and AS method on two datasets.
+Table 3. The results of swapping test sets based on Patch method and AS method on two datasets.
 Training Sets Method Test Sets Results (×$(\times 10^{-3})$
@@
 Architecture Impact: This article breaks down the important modules of PU-GCN, aiming to explore which part has the greatest impact on point cloud upsampling.
-To this end, the original network backbone has been increased or decreased in three ways: removing DenseGCN (b), adding Refiner (c), and removing DenseGCN while adding Refiner (d), all of them are showed in Fig.
-3.
+To this end, the original network backbone has been increased or decreased in three ways: removing DenseGCN (b), adding Refiner (c), and removing DenseGCN while adding Refiner (d), all of them are showed in Figure  2.
 The experiment was trained on the PU1K dataset and tested on the test sets of the PU1K dataset and ABC10K dataset.
-The results are shown in Tab.
-4.
+The results are shown in Table  4.
 The results showed that upsampling on the test set with three resolutions of 256-1024, 512-2048 and 2048-8192 in PU1K performed better than the original PU-GCN with DenseGCN, regardless of the presence or absence of Refiner, under the condition of removing DenseGCN.
 Under the upsampling scale of 2048-8192, the uniform segmentation method that removes DenseGCN is better than the Patch method.
-Table 4: Perform ablation experiments on different architectures of PU-GCN using two data input methods.
+Table 4. Perform ablation experiments on different architectures of PU-GCN using two data input methods.
 Network Backbone Method Test Sets Results (×$(\times 10^{-3})$ Model size (MB)
@@
 AS 14.961 27.822
-without Inception DenseGCN Patch PU1K $256\to 1024$ 35.061 54.340 99.152
+w/o Inception DenseGCN Patch PU1K $256\to 1024$ 35.061 54.340 99.152
 AS 35.229 52.819
@@
 AS 13.504 23.130
-with Refiner Patch PU1K $256\to 1024$ 35.075 54.227 87.773
+w/ Refiner Patch PU1K $256\to 1024$ 35.075 54.227 87.773
 AS 35.860 53.652
@@
 AS 14.817 26.860
-with Refiner but without Inception DenseGCN Patch PU1K $256\to 1024$ 34.309 54.070 130.945
+w/ Refiner w/o Inception DenseGCN Patch PU1K $256\to 1024$ 34.309 54.070 130.945
 AS 35.158 53.434
@@
 AS 13.441 22.899
-3.5 Visualizations
-Visualize the above experimental results now, which directly displays qualitative comparisons.
-The qualitative results of upsampling using different point cloud datasets are shown in Fig.
-4 and Fig.
-5.
+3.5. Visualizations
+The qualitative results of upsampling using different point cloud datasets are shown in Figure  3 and Figure  4.
 It can be seen that the input based on Patch performs better in distinguishing edges on the boundaries compared to the AS method.
-This is particularly evident on the CAD based ABC dataset, as shown in Fig.
-5.
-However, in Fig.
-4, it can be seen that the AS method performs better locally in the model.
+This is particularly evident on the CAD based ABC dataset, as shown in Figure  4.
+However, in Figure  3, it can be seen that the AS method performs better locally in the model.
 Relatively speaking, it is denser and the connections between points are tighter.
 In addition, even though DenseGCN performs better on CD and HD on the PU1K dataset, this results in the point cloud lacking local features in visualization.
-On the other hand, Refiner can solve the problem of identifying local features in the model, achieving the utilization of both global and local features.
-Figure 4: Visualization of PU-GCN with Different Architectures on the PU1K Dataset.
-Figure 5: Visualization example on the ABC dataset.
-4 Discussion and Future Work
-4.1 Discussion
+Figure 3. Visualization of PU-GCN with Different Architectures on the PU1K Dataset.
+Figure 4. Visualization example on the ABC dataset.
+4. Discussion
 Now, we will focus on why the patch upsampling method performs better in PU-GCN experiments.
@@
 This phenomenon can be explained by the following points:
-More effective local feature extraction: When upsampling on AS method, each AS slice contains fewer point clouds, and the graph convolutional network can extract local features more finely.
-Due to each AS slice of the point cloud being within a small range, local features are easier to capture and reconstruct, resulting in excellent performance on CD and HD.
-Set $\mathcal{P}_{k}$ represents the $k$-th AS slice, $\mathcal{P}_{k}=\{\mathbf{p}_{i}\mid i\in I_{k}\}$, where $I_{k}$ is the point index set for AS $\mathcal{P}_{k}$.
-For each AS slice $\mathcal{P}_{k}$, its point cloud can be represented as $\mathbf{X}_{k}\in\mathbb{R}^{N_{k}\times d}$, where $N_{k}$ is the number of AS slices and $d$ is the dimension of the points.
-In a graph convolutional network, the output of the $l$-th layer is represented as:
-$\mathbf{H}_{k}^{(l)}=\sigma(\mathbf{A}_{k}\mathbf{H}_{k}^{(l-1)}\mathbf{W}^{(l)})$
-In an AS slice, the adjacency matrix $\mathbf{A}_{k}$ is denser, with tighter connections between points, making convolution operations more effective in capturing local features.
-Local consistency of AS slices: When dealing with smaller AS slices, the local consistency of point clouds is stronger, that is, the similarity between points is higher.
-This enables graph convolutional networks to better utilize these local consistency for learning and reconstruction, reducing local errors introduced by Patch based method.
-The AS method deals with point clouds with strong local consistency, which can be measured by the local covariance matrix.
-For AS slice $\mathcal{P}_{k}$, its point cloud $\mathbf{X}_{k}$ have the local covariance matrix is:
-$\mathbf{C}_{k}=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}(\mathbf{x}_{i}-\mathbf{\mu}_{k})(\mathbf{x}_{i}-\mathbf{\mu}_{k})^{\top}$
-Where $\mathbf{\mu}_{k}$ is mean vector of midpoint of AS slice $\mathcal{P}_{k}$ :
-$\mathbf{\mu}_{k}=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}\mathbf{x}_{i}$
-Local covariance matrix $\mathbf{C}_{k}$ can measure the local consistency of point clouds.
-In the case of strong local consistency, the smaller feature values of $\mathbf{C}_{k}$ make local feature extraction more effective.
 The effectiveness of boundary processing strategies: Although theoretically AS sampling followed by segmentation can avoid boundary problems, in practical applications, the overlapping and smoothing of boundaries in patch segmentation may be more effective than those after AS sampling, thereby improving the overall quality of upsampling.
@@
 $\mathbf{X}_{k,m}^{\prime}=\alpha\mathbf{X}_{k,m}^{(k)}+(1-\alpha)\mathbf{X}_{k,m}^{(m)}$
-Among them:
+where:
   • $\mathbf{X}_{k,m}^{(k)}$ and $\mathbf{X}_{k,m}^{(m)}$ respectively represent overlapping area point clouds in Patch $\mathcal{P}_{k}$ and $\mathcal{P}_{m}$.
@@
 Due to the segmented subset $\mathcal{P}_{k}$ lack of local consistency may lead to a decrease in the quality of feature extraction and reconstruction.
-4.2 Future Work
-In point cloud 3D model sampling, such as Pu-Edgeformr++ [10], it can achieve good point cloud edge recognition and sampling.
-However, taking the PU-GCN model used in this article as an example, the existing point cloud upsampling models are not convincing in their performance on the ABC dataset.
-The method of reverse engineering CAD models follows a series of typical steps: data capture and preprocessing, point cloud segmentation, face classification, and analysis of faces generated by segmentation methods [3].
-The poor quantitative and qualitative performance on the CAD dataset ABC may be some challenges due to the heterogeneous and unstructured nature of manufacturing data [16].
-In the future, the application of point cloud upsampling in reverse engineering and CAD datasets may require a similar approach to Point2CAD [14], which involves dividing point clouds into clusters corresponding to CAD model faces, fitting each cluster with an analytical surface primitive, and proposing a novel fitting scheme for free-form surfaces based on neural representation.
-5 Conclusion
-In order to ensure the model integrity of point cloud upsampling, this article mainly studies the input of point cloud data.
+5. Future Works
+PU-GCN model used in this article as an example, the existing Patch based upsampling method is better than AS method we proposed.
+This may be due to the difference in dataset size between the two methods.
+In future work, we will delve deeper into the differences in interpretability between Patch and AS methods through representation learning methods.
+At the same time, due to the similarity between the use of point clouds and the completion of point clouds.
+We will refer to the method in article  (Xia et al., 2021) and use encoders for the data input of Patch and AS respectively, in order to merge global and edge features and achieve better results.
+6. Conclusion
+To ensure the model integrity of point cloud upsampling, this article mainly studies the input of point cloud data.
 Specifically, this article proposes a new data input method, also known as a new training model method, called Average Segment.
@@
 The traditional patch method still performs slightly better in both CD and HD indicators, but through relevant ablation research and visualization, we can see that the AS method can maintain local features within the model.
-In future work, we consider combining point cloud models with mesh face clusters to perform pairwise surface intersections to identify object edges, thereby achieving better edge, local, and global representations.

### v3

@@
 Rethinking Data Input for Point Cloud Upsampling
-Abstract.
+Abstract
 Point cloud upsampling is crucial for tasks like 3D reconstruction.
@@
 To understand this, we will delve into factors in feature extraction, and network architecture that influence upsampling results.
+…
+keywords:
 Deep learning, Point cloud, Upsampling, Data input, 3D reconstruction
-ccs: Computing methodologies Computer vision
-1. Introduction
+1 Introduction
 The improvement of 3D scanners in 3D sensors, as well as the rapid development of deep learning in computer vision, have led to a gradual shift in the popularity of computer vision technology from 2D image to 3D model.
-Especially in the direction of autonomous driving and robots that exchange with the physical world, there is a particular focus on three-dimensional vision (Endres et al., 2013; Roriz et al., 2021; Naseer et al., 2018; Zhou and Tuzel, 2018).
-Figure 1. The original Mesh models have been transformed into a point cloud through patch based sampling and the Average Segment method proposed in this article.
+Especially in the direction of autonomous driving and robots that exchange with the physical world, there is a particular focus on three-dimensional vision [1, 2, 3, 4].
+Figure 1: The original Mesh models have been transformed into a point cloud through patch based sampling and the Average Segment method proposed in this article.
 Vis-a-vis other representations such as Mesh, point clouds are an effective way to represent three-dimensional shapes that are widely used in the aforementioned fields.
-In addition, the task of reverse engineering CAD models from 3D point clouds is receiving increasing attention (Alai, 2013).
-In the process of redesigning the model obtained from reverse engineering, although there are Application Programming Interface (API) available in CAD software (Kao and Lin, 1998), if these designs involve reverse engineering and highly customized and personalized designs, CAD APIs are powerless (Shatnawi et al., 2017).
+In addition, the task of reverse engineering CAD models from 3D point clouds is receiving increasing attention [5].
+In the process of redesigning the model obtained from reverse engineering, although there are Application Programming Interface (API) available in CAD software [6], if these designs involve reverse engineering and highly customized and personalized designs, CAD APIs are powerless [7].
 Therefore, the generation of model point clouds is crucial.
-The point cloud upsampling technology based on deep learning has developed with the development of deep learning, extracting multi-resolution features from the initial PU-Net (Yu et al., 2018) on the basis of pointnet++ (Qi et al., 2017).
+The point cloud upsampling technology based on deep learning has developed with the development of deep learning, extracting multi-resolution features from the initial PU-Net [8] on the basis of pointnet++ [9].
 Then, multiple independent multi-layer perceptrons (MLPs) are introduced to generate the required dense point cloud.
-Moving on to MPU/3PU (Yifan et al., 2019), KNN is used for interpolation and EdgeConv (Wang et al., 2019) is used to extract features for hierarchical reconstruction through recursive upsampling.
-Implementing upsampling PU-GAN (Li et al., 2019) based on GAN (Goodfellow et al., 2020) generation architecture, guiding the generator to upsample and approximate the real point cloud.
-PU-GCN (Qian et al., 2021) modifies the upsampling module from simple duplicate to Nodeshuffle based on feature extraction in MPU, achieving very good results.
-And Transformer’s point cloud upsampling network (Qiu et al., 2022; Vaswani et al., 2017).
-There are also methods to create coarse point clouds and then refine them by combining global and local methods (Li et al., 2021).
+Moving on to MPU/3PU [10], KNN is used for interpolation and EdgeConv [11] is used to extract features for hierarchical reconstruction through recursive upsampling.
+Implementing upsampling PU-GAN [12] based on GAN [13] generation architecture, guiding the generator to upsample and approximate the real point cloud.
+PU-GCN [14] modifies the upsampling module from simple duplicate to Nodeshuffle based on feature extraction in MPU, achieving very good results.
+And Transformer’s point cloud upsampling network [15, 16].
+There are also methods to create coarse point clouds and then refine them by combining global and local methods [17].
 The above methods mainly utilize Euclidean spatial features at potential spatial points, upsampling from low resolution to high resolution.
-The inherent limitations have also been addressed by the work of upsampling point clouds at any scale (Zhao et al., 2022; He et al., 2023) and capturing local information (Kim et al., 2024).
+The inherent limitations have also been addressed by the work of upsampling point clouds at any scale [18, 19] and capturing local information [20].
 Nevertheless, these works still have limitations, which are on the training dataset.
-Existing point cloud upsampling work (Qian et al., 2021; Qiu et al., 2022; Zhao et al., 2022; He et al., 2023; Kim et al., 2024; Mao et al., 2022) without exception used the PU1K dataset (Qian et al., 2021), which was used based on the methods of MPU and PU-GAN (Yifan et al., 2019; Li et al., 2019).
+Existing point cloud upsampling work [14, 15, 18, 19, 20, 21] without exception used the PU1K dataset [14], which was used based on the methods of MPU and PU-GAN [10, 12].
 To obtain training data by randomly sampling the entire 3D model using Patch under conditions of limited computing power, especially limited graphics memory.
@@
 Moreover, the existing point cloud upsampling work focuses on 3D models similar to the PU1K dataset, while neglecting the exploration of point cloud upsampling for parts and CAD.
-In addition to conducting experiments on the PU1K dataset, this article will also discuss the upsampling effect of point cloud models on workpieces on the ABC dataset (Koch et al., 2019).
+In addition to conducting experiments on the PU1K dataset, this article will also discuss the upsampling effect of point cloud models on workpieces on the ABC dataset [22].
 Based on the above issues, in order to explore why numerous point cloud sampling data are still based on Patch input, this paper proposes a new data input method.
 Unlike dividing point clouds according to Patch, this article takes the point cloud model uniformly and randomly divided and corresponds them one by one as input, which is named as Average Segment (AS).
-After verifying the method proposed in this article, experiments were conducted on point cloud upsampling using the PU-GCN (Qian et al., 2021) network, and cross validation was performed on both the PU1k and ABC datasets.
-And during the experiment, the encoder and decoder of PU-GCN (Qian et al., 2021) were ablated.
+After verifying the method proposed in this article, experiments were conducted on point cloud upsampling using the PU-GCN [14] network, and cross validation was performed on both the PU1k and ABC datasets.
+And during the experiment, the encoder and decoder of PU-GCN [14] were ablated.
 To solve the above problems, this paper proposes a new data input method.
 Unlike dividing point clouds according to patches, this paper uniformly and randomly divides point cloud models and uses them as inputs one by one.
-After verifying the method proposed in this article, experiments were conducted on point cloud upsampling using the PU-GCN (Qian et al., 2021), and cross validation was performed on both the PU1k (Qian et al., 2021) and ABC datasets (Koch et al., 2019).
-And during the experiment, the encoder and decoder of PU-GCN (Qian et al., 2021) were ablated.
+After verifying the method proposed in this article, experiments were conducted on point cloud upsampling using the PU-GCN [14], and cross validation was performed on both the PU1k [14] and ABC datasets [22].
+And during the experiment, the encoder and decoder of PU-GCN [14] were ablated.
 The contributions of this article are as follows:
@@
 -Conducting ablation experiments on the main structure of PU-GCN, determining that upsampling mainly relies on the extended feature modules in the decoder, and pointing out the key link in point cloud generation.
-2. Methodology
-2.1. Overview
+2 Methodology
+2.1 Overview
 As shown in Figure  1, given a sparse point cloud $\mathcal{P}\in\mathbb{R}^{N\times 3}$, the input methods based on Patch and the AS method proposed in this paper are used as data inputs, with PU-GCN as the network backbone, to generate a dense point cloud $\mathcal{P^{\prime}}\in\mathbb{R}^{rN\times 3}$, where $r$ is the upsampling scale.
@@
 Although the overall local features are retained during sampling, point cloud data may not be continuous enough in some areas due to random sampling, and further processing may be needed to ensure the continuity and consistency of the point cloud.
-2.2. Data Input Based on Patch
+2.2 Data Input Based on Patch
 In this case, it is necessary to divide the Mesh into multiple patches.
@@
 $\mathcal{P}^{\prime}=\bigcup_{k=1}^{K}\mathcal{P}_{k}^{\prime}$
-2.3. Data Input Based on Average Segment
+2.3 Data Input Based on Average Segment
 We can assume that the distribution of sampling points follows a uniform distribution.
@@
 Algorithm 1 Point Cloud Upsampling with Patch Segmentation
-3. Experiments
-3.1. Settings
+3 Experiments
+3.1 Settings
 Datasets: In general speaking, the experiment in this article applied the following two 3D benchmark datasets:
-  • PU1K: This is a point cloud upsampling dataset introduced in PU-GCN (Qian et al., 2021).
-Overall, PU1K consists of 1147 3D models, including 120 3D models from the PU-GAN dataset, as well as 900 different models collected from ShapeNetCore (Chang et al., 2015).
+  • PU1K: This is a point cloud upsampling dataset introduced in PU-GCN [14].
+Overall, PU1K consists of 1147 3D models, including 120 3D models from the PU-GAN dataset, as well as 900 different models collected from ShapeNetCore [23].
 The test set includes 27 models from PU-GAN and over 100 models from ShapeNetCore.
 Covering 50 object categories.
-  • ABC Dataset: The ABC Dataset (Koch et al., 2019) is a dataset containing one million computer-aided design (CAD) models, aimed at promoting research on geometric deep learning methods and applications.
+  • ABC Dataset: The ABC Dataset [22] is a dataset containing one million computer-aided design (CAD) models, aimed at promoting research on geometric deep learning methods and applications.
 It is currently one of the largest publicly available datasets for geometric deep learning research.
@@
 Notwithstanding, a batch size of $32$ with $50$ training epochs was used, with an initial learning rate of $5\times 10^{-4}$, a decay rate of $0.05$, and so on.
-Figure 2.
-The PU-GCN network structure used in the experiment is shown in the figure, where (a) is the original PU-GCN structure.
+Figure 2: The PU-GCN network structure used in the experiment is shown in the figure, where (a) is the original PU-GCN structure.
 (b), (c) and (d) are all architectures used in the ablation experiment to explore the modules that have the greatest impact on upsampling.
-3.2. Evaluation Metrics
-During the testing process, this article adopted the common practices used in previous work on point cloud upsampling (Qian et al., 2021; Qiu et al., 2022; Zhao et al., 2022; He et al., 2023; Kim et al., 2024; Mao et al., 2022).
+3.2 Evaluation Metrics
+During the testing process, this article adopted the common practices used in previous work on point cloud upsampling [14, 15, 18, 19, 20, 21].
 That is to say, this article will cut the input point cloud into multiple subset Patch and AS of all N points.
@@
 Each test sample has a low resolution point cloud consisting of 256, 512 and 1048 points, and a high-resolution point cloud consisting of 1024,2048 and 8196 points.
-This article quantitatively evaluates the upsampling performance of PU-GCN under two methods based on three widely used indicators: (i) Chamfer distance (CD) (Hajdu et al., 2012), (ii) Hausdorff distance (HD) (Berger et al., 2013), and (iii) model parameter quantity.
-3.3. Results
+This article quantitatively evaluates the upsampling performance of PU-GCN under two methods based on three widely used indicators: (i) Chamfer distance (CD) [24], (ii) Hausdorff distance (HD) [25], and (iii) model parameter quantity.
+3.3 Results
 PU1K dataset: Table  1 shows the quantitative results of using patch method and uniform segmentation method with PU-GCN network training on the PU1K dataset.
 By upsampling at resolutions of 256-1024 and 2048-8192, it can be seen that the Patch based method outperforms the AS method proposed in this paper in all two metrics of distance measurement.
-Table 1. The data input based on the Patch method and AS method is obtained using PU-GCN in PU1K.
+Table 1: The data input based on the Patch method and AS method is obtained using PU-GCN in PU1K.
 Datasets Test Scale Method Results (×$(\times 10^{-3})$
@@
 As shown in Table  2, in the 4$\times$ upsampling experiment, the two evaluation metrics of the Patch based method outperformed the AS method proposed in this paper in all two metrics of distance measurement.
-Table 2. The data input based on the Patch method and AS method is obtained using PU-GCN in ABC10K.
+Table 2: The data input based on the Patch method and AS method is obtained using PU-GCN in ABC10K.
 Datasets Test Scale Method Results (×$(\times 10^{-3})$
@@
 Ergo, local features are easier to capture and reconstruct, resulting in excellent performance on CD and HD.
-3.4. Ablations
+3.4 Ablations
 Test sets cross validation: This article will conduct cross validation on the PU-GCN trained on the PU1K and ABC10K datasets under the validation sets of both datasets.
@@
 However, when trained on the ABC10K dataset and tested on the PU1K dataset, there was no significant difference between the CD and HD indicators.
-Table 3. The results of swapping test sets based on Patch method and AS method on two datasets.
+Table 3: The results of swapping test sets based on Patch method and AS method on two datasets.
 Training Sets Method Test Sets Results (×$(\times 10^{-3})$
@@
 Under the upsampling scale of 2048-8192, the uniform segmentation method that removes DenseGCN is better than the Patch method.
-Table 4. Perform ablation experiments on different architectures of PU-GCN using two data input methods.
+Table 4: Perform ablation experiments on different architectures of PU-GCN using two data input methods.
 Network Backbone Method Test Sets Results (×$(\times 10^{-3})$ Model size (MB)
@@
 AS 13.441 22.899
-3.5. Visualizations
+3.5 Visualizations
 The qualitative results of upsampling using different point cloud datasets are shown in Figure  3 and Figure  4.
@@
 In addition, even though DenseGCN performs better on CD and HD on the PU1K dataset, this results in the point cloud lacking local features in visualization.
-Figure 3. Visualization of PU-GCN with Different Architectures on the PU1K Dataset.
-Figure 4. Visualization example on the ABC dataset.
-4. Discussion
+Figure 3: Visualization of PU-GCN with Different Architectures on the PU1K Dataset.
+Figure 4: Visualization example on the ABC dataset.
+4 Discussion
 Now, we will focus on why the patch upsampling method performs better in PU-GCN experiments.
@@
 Due to the segmented subset $\mathcal{P}_{k}$ lack of local consistency may lead to a decrease in the quality of feature extraction and reconstruction.
-5. Future Works
+5 Future Works
 PU-GCN model used in this article as an example, the existing Patch based upsampling method is better than AS method we proposed.
@@
 At the same time, due to the similarity between the use of point clouds and the completion of point clouds.
-We will refer to the method in article  (Xia et al., 2021) and use encoders for the data input of Patch and AS respectively, in order to merge global and edge features and achieve better results.
-6. Conclusion
+We will refer to the method in article  [26] and use encoders for the data input of Patch and AS respectively, in order to merge global and edge features and achieve better results.
+6 Conclusion
 To ensure the model integrity of point cloud upsampling, this article mainly studies the input of point cloud data.